# Perceptual Decoupling for Scalable Multi-modal Reasoning via Reward-Optimized Captioning

### v1

@@ C.2 Details for Pairwise Comparisons
We present the details of the human evaluation conducted for the pairwise comparison experiment. For this, 100 questions are randomly sampled from the testmini set of MathVista, and captions are generated using Qwen2.5-VL-3B, trained with and without CRO. A total of 4 trained human annotators are recruited, with each annotator responsible for evaluating 25 questions to determine a winner or a tie.
The annotators’ primary task is to evaluate a question and its two anonymous model-generated answers, then select the one they deem of higher quality, following the scoring principle adopted from [4]. If neither answer stood out, they are instructed to choose "tie."
In Table 6, we report the win/tie/lose ratio (i.e., “win” means captions generated by MLLMs with CRO is better) and an additional measure of consistency, calculated by the agreement rate between GPT-4o and human judgments. As can be seen, Qwen2.5-VL-3B trained with CRO demonstrates superior caption quality under human evaluation. This aligns with the result in Figure 7, which is further supported by the high consistency of 85%. This supports the rationale of using GPT-4o as a judge for evaluating caption quality.
@@
Win Tie Lose Consistency
60% 29% 11% 85%
C.3 Training Dynamics of CRO
We show the average reward scores and caption lengths over training steps in Figure 13. We observe the following:
  • Rewards increased as training progresses. This confirms the effectiveness of CRO as it allows the MLLM to generate captions that lead to higher reasoning accuracy.
  • Caption lengths grow as training progresses. An explanation for this phenomenon is that the MLLM learns to generate more comprehensive captions during training, which is reflected by longer lengths. This is also confirmed in the Appendix D, where we visualize the captions.
@@ C.3 Training Dynamics of CRO
Figure 13: Training dynamics of the CRO process: Reward and Length Over Steps for Qwen2.5-VL-3B, Qwen2.5-VL-7B and Qwen2.5-VL-32B
Appendix D Case Study
@@
We conduct a case study on the generated query-conditioned captions. Specifically, for a multi-modal reasoning question and image, we investigate the quality of the generated captions. For MLLMs, we consider Qwen2.5-VL series (3B/32B) both with and without CRO. We visualize the question, image and captions in Tables LABEL:tbl:case_3b_1- LABEL:tbl:case_3b_7 for Qwen2.5-VL-3B and Tables LABEL:tbl:case_32b_1- LABEL:tbl:case_32b_5 for Qwen2.5-VL-32B.
Comparing the captions generated by MLLMs with and without CRO, we discover the following:
  • CRO leads to more visual details. We highlight these visual details in red in the table. Notably, these details are important clues required to correctly solve the question. This shows that CRO is effective in improving the quality (especially comprehensiveness) of the query-conditioned captions.
  • CRO leads to captions with more organized and hierarchical structures. For example, in Table LABEL:tbl:case_3b_6, the MLLM with CRO describes the images at three levels, i.e., Tropic level, Terrestrial food chain and aquatic food chain. This allows the reasoner to quickly locate important information in the captions. However, the original MLLM uses a sequence of sentences that are less clear.
@@
In Table LABEL:tbl:r_case_3b_1, we provide a complete comparison of captions (generated by Qwen2.5-VL-3B with and without CRO) and the resulting reasoning results (produced by R1-7B). Similarly to the case study on caption quality, MLLMs trained with CRO generate captions that capture more details. This is critical to the correctness of the reasoning process. As can be seen, the reasoner that receives captions with CRO arrives at the correct answer after several rounds of thinking and reflection. However, the reasoner that accepts the under-optimized caption experiences multiple contradictions and confusion (highlighted in brown), which leads to responses that exceeds the maximum context length and finally fails this problem.
Table 7: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-3B):
@@
Table 8: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-3B):
@@
Table 9: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-3B):
@@
Table 10: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-3B):
@@
Table 11: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-3B):
@@
Table 12: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-3B): The image shows a food chain diagram with two types of food chains: terrestrial and aquatic. The terrestrial food chain includes a caterpillar, lizard, fish, and snake, with the caterpillar being the primary producer and the snake being the tertiary consumer. The aquatic food chain includes zooplankton, phytoplankton, fish, and a pelican, with the zooplankton being the primary producer and the pelican being the tertiary consumer.
Table 13: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-3B): The image is a food web diagram depicting various organisms and their interactions within an ecosystem. It includes a deer, trees, grasses, mice, crickets, frogs, hawks, and owls. The arrows indicate the direction of energy flow from producers (trees and grasses) to consumers (deer, mice, crickets, frogs, hawks, and owls), and decomposers (decomposers). The diagram shows how energy moves through the ecosystem, with each organism feeding on others, and how the removal of frogs could potentially lead to an increase in the population of crickets, as they are a primary food source for frogs.
Table 14: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-32B):
@@
Table 15: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-32B):
@@
Table 16: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-32B):
@@
Table 17: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-32B):
@@
Table 18: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
@@
Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-32B):
[…]
Table 19: Reasoning Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red. Reasoning process that show confusing and contradiction are highlighted in brown.
@@
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
@@
Query-conditioned Captions (Qwen2.5-VL-3B):
@@
Reasoning Result of R1-7B (captions from Qwen2.5-VL-3B w/ CRO):
Alright, let me try to figure out this geometry problem step by step. So, we have triangle $ABC$ inscribed in a circle with center $O$. $AB$ and $BC$ are both 16 units long, making it an isosceles triangle with $AB=BC$. The diagram shows that when arc $AB$ is folded along chord $AB$, it intersects $BC$ at point $D$, which is the midpoint of $BC$. So, $BD=DC=8$ units.
@@
Appendix E Broader Impacts
RACRO improves multi-modal reasoning by aligning perception with reasoning through reinforcement learning. While beneficial for tasks requiring precise vision-language understanding, this alignment could be misused for generating misleading content or surveillance. Caution is advised when applying RACRO in sensitive domains. Mitigation strategies include human oversight, responsible deployment, and transparency in model behavior.
Appendix F Limitation
We evaluate the effectiveness of the proposed RACRO framework exclusively on multi-modal reasoning tasks. However, its potential applicability could extend to other important domains, such as medical visual question answering (VQA) and geospatial reasoning. Exploring these areas represents promising directions for future work.
Appendix G More Discussions
Future works.
This work aims at decoupling multi-modal reasoning with visual perception and LLM reasoning, and the successful visual perception is essential for downstream reasoning. How to further enhance the visual perception capability is a prevailing research direction (e.g., utilizing AIGC-generated data [10, 16, 17, 18, 23, 28, 29, 33, 50] and enhanced visual representations [7, 8, 35, 58])

### v2

@@ E.3 Details for Pairwise Comparisons
We present the details of the human evaluation conducted for the pairwise comparison experiment. For this, 100 questions are randomly sampled from the testmini set of MathVista, and captions are generated using Qwen2.5-VL-3B, trained with and without VPO. A total of 4 trained human annotators are recruited, with each annotator comparing all the captions pairs to determine a winner or a tie. For each caption pair, we aggregate the results from different annotators by taking the majority of the decisions. Specifically, there are 4 annotators and only 3 decisions (win, tie and lose), so there is at least one decision that occurs twice. We compute the inter-annotator consistency following Zheng et al. (2023) by calculating the ratio of identical decision pairs out of all possible decision pairs and average them across all samples.
In Table 8, we report the win/tie/lose ratio (i.e., “win” means captions generated by MLLMs with VPO is better) and an additional measure of GPT-human consistency, calculated by the agreement rate between GPT-4o and human judgments. As can be seen, Qwen2.5-VL-3B trained with VPO demonstrates superior caption quality under human evaluation. This aligns with the result in Figure 11, which is further supported by the high consistency of 87%. This supports the rationale of using GPT-4o as a judge for evaluating caption quality.
@@
Win Tie Lose GPT-human consistency Inter-annotator consistency
62% 32% 6% 87% 85%
E.4 Training Dynamics of VPO
We show the average reward and caption lengths over training in Figures 21 and 22. We observe that:
  • Rewards increased as training progresses. This confirms the effectiveness of VPO as it allows the MLLM to generate captions that lead to higher reasoning accuracy.
  • Caption lengths grow as training progresses. An explanation for this phenomenon is that the MLLM learns to generate more comprehensive captions during training, which is reflected by longer lengths. This is also confirmed in the Appendix G, where we visualize the captions.
@@ E.4 Training Dynamics of VPO
(d) Qwen2.5-VL-72B
Figure 21: Rewards over training steps for VPO.
(a) Qwen2.5-VL-3B
(b) Qwen2.5-VL-7B
(c) Qwen2.5-VL-32B
(d) Qwen2.5-VL-72B
Figure 22: Caption length over training steps for VPO.
E.5 Training Dynamics of GRPO
We visualize the training dynamics of GRPO in Figure 23. For 3B and 7B MLLMs, the rewards fluctuate and eventually drop after an initial convergence. In contrast, the 72B model exhibits a more stable convergence. In both cases, subsequent VPO stage continues to improve the performance under the perception-reasoning pipelines.
(a) Qwen2.5-VL-3B
(b) Qwen2.5-VL-7B
(c) Qwen2.5-VL-72B
Figure 23: Rewards over training steps for GRPO.
Appendix F Use of Large Language Models (LLMs)
In this paper, the role of Large Language Models (LLMs) was confined to a minor, supporting capacity for polishing the writing. They were not involved in the core research process, such as ideation or analysis.
Appendix G Case Study
@@
We conduct a case study on the generated query-relevant captions. Specifically, for a multi-modal reasoning question and image, we investigate the quality of the generated captions. For MLLMs, we consider Qwen2.5-VL series (3B/32B) both with and without VPO. We visualize the question, image and captions in Tables LABEL:tbl:case_3b_1- LABEL:tbl:case_3b_7 for Qwen2.5-VL-3B and Tables LABEL:tbl:case_32b_1- LABEL:tbl:case_32b_5 for Qwen2.5-VL-32B.
Comparing the captions generated by MLLMs with and without VPO, we discover the following:
  • VPO leads to more visual details. We highlight these visual details in red in the table. Notably, these details are important clues required to correctly solve the question. This shows that VPO is effective in improving the quality (especially comprehensiveness) of the query-relevant captions.
  • VPO leads to captions with more organized and hierarchical structures. For example, in Table LABEL:tbl:case_3b_6, the MLLM with VPO describes the images at three levels, i.e., Tropic level, Terrestrial food chain and aquatic food chain. This allows the reasoner to quickly locate important information in the captions. However, the original MLLM uses a sequence of sentences that are less clear.
@@
In Table LABEL:tbl:r_case_3b_1, we provide a complete comparison of captions (generated by Qwen2.5-VL-3B with and without VPO) and the resulting reasoning results (produced by R1-7B). Similarly to the case study on caption quality, MLLMs trained with VPO generate captions that capture more details. This is critical to the correctness of the reasoning process. As can be seen, the reasoner that receives captions with VPO arrives at the correct answer after several rounds of thinking and reflection. However, the reasoner that accepts the under-optimized caption experiences multiple contradictions and confusion (highlighted in brown), which leads to responses that exceeds the maximum context length and finally fails this problem.
Appendix H Limitation
Auto-thinking.
RAPID is specifically designed for multi-modal reasoning, and it is appealing to explore how to flexibly switch between fast and slow thinking dependent on the complexity of input queries, without human prior and prompt engineering.
Domain-Specific Design.
The RAPID architecture, in its current form, is optimized for multi-modal math and science reasoning. Its extension to other domains, such as spatial reasoning (Wang et al., 2024a), would require more than re-evaluation; it would demand specific adaptations to the model itself. The experiments on general benchmarks reported in Sec. 4.1 do not test the full reasoning pipeline, as the LLM was not activated. Therefore, investigating the adaptations required to generalize RAPID remains a key open direction for future research.
Adapting the LLM.
In the current implementation of RAPID, the LLM functions as a static reasoning module with its parameters kept frozen. A valuable direction for future work is to explore adapting the LLM to our perception-reasoning framework, potentially through methods like supervised fine-tuning or reinforcement learning. While this would introduce additional computational overhead for training, it remains an open empirical question whether the potential performance gains would justify the increased cost.
Table 9: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@ Appendix H Limitation
Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-3B):
@@
Table 10: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-3B):
@@
Table 11: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-3B):
@@
Table 12: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-3B):
@@
Table 13: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-3B):
@@
Table 14: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-3B): The image shows a food chain diagram with two types of food chains: terrestrial and aquatic. The terrestrial food chain includes a caterpillar, lizard, fish, and snake, with the caterpillar being the primary producer and the snake being the tertiary consumer. The aquatic food chain includes zooplankton, phytoplankton, fish, and a pelican, with the zooplankton being the primary producer and the pelican being the tertiary consumer.
Table 15: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-3B): The image is a food web diagram depicting various organisms and their interactions within an ecosystem. It includes a deer, trees, grasses, mice, crickets, frogs, hawks, and owls. The arrows indicate the direction of energy flow from producers (trees and grasses) to consumers (deer, mice, crickets, frogs, hawks, and owls), and decomposers (decomposers). The diagram shows how energy moves through the ecosystem, with each organism feeding on others, and how the removal of frogs could potentially lead to an increase in the population of crickets, as they are a primary food source for frogs.
Table 16: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-32B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-32B):
@@
Table 17: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-32B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-32B):
@@
Table 18: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-32B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-32B):
@@
Table 19: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-32B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-32B):
@@
Table 20: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
@@
Query-relevant Captions (Qwen2.5-VL-32B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-32B):
The image shows five groups of cartoon-style cats, each labeled with a letter (A, B, C, D, E). Each group contains multiple cats with various patterns and colors, including striped, spotty, and solid-colored cats. The cats are depicted in different poses, such as sitting, lying down, or standing. The question asks to identify the group that matches the description of Misty’s kittens, which includes two striped, one spotty, and two white kittens, with one having differently colored ears. Detailed Description of Each Group: Group A: Contains cats with striped and spotty patterns. Some cats have differently colored ears. Group B: Features cats with striped and solid-colored patterns. No cats have differently colored ears. Group C: Includes cats with striped, spotty, and solid-colored patterns. No cats have differently colored ears. Group D: Shows cats with striped and solid-colored patterns. No cats have differently colored ears. Group E: Contains cats with striped, spotty, and solid-colored patterns. No cats have differently colored ears. The task is to identify the group that matches the specific description of Misty’s kittens. Misty’s Kittens: Two striped kittens One spotty kitten Two white (solid-colored) kittens One kitten has differently colored ears Conclusion: Only Group A meets all the criteria, including the presence of differently colored ears. $\boxed{\text{A}}$
Table 21: Reasoning Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red. Reasoning process that show confusing and contradiction are highlighted in brown.
@@
Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
@@
Query-relevant Captions (Qwen2.5-VL-3B):
@@
Reasoning Result of R1-7B (captions from Qwen2.5-VL-3B w/ VPO):
Alright, let me try to figure out this geometry problem step by step. So, we have triangle $ABC$ inscribed in a circle with center $O$. ABAB and $BC$ are both 16 units long, making it an isosceles triangle with $AB=BC$. The diagram shows that when arc $AB$ is folded along chord $AB$, it intersects $BC$ at point $D$, which is the midpoint of $BC$. So, $BD=DC=8$ units.